OpenStudy (anonymous):

Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + 5y inches, the other side should be (x ^ 2 -5xy - 4y^ 2) inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution.

OpenStudy (anonymous):

Area is Length X Breadth area = (x+5y)(x^2-5xy-4y^2) = x^3 - 5yx^2 - 4xy^2 + 5yx^2 - 25xy^2 - 20y^3 = x^3 - 29xy^2 - 20y^3 (SQUARE INCHES)